Find the exact values of sin 2u , Cos 2u, and tan 2u using double angle formulas.
1. csc u= 3, pi/2 < u > pi

Answer by stanbon(75887)

Find the exact values of sin 2u , Cos 2u, and tan 2u using double angle formulas.
1. csc u= 3, pi/2 < u > pi; QII where x is negative and y is positive.
sin(u) = 1/3; y = 1 and r = 3
----------------------------------------
x = -sqrt[9-1] = -sqrt(8)
cos(u) = -sqrt(8)/3 = -(2/3)sqrt(2)
---------------------------------------
tan(u) = y/x = -1/sqrt(8)
===========
Your Problems:
sin(2u) = 2sin(u)cos(u) = 2[1/3][-sqrt(8)/3]
---
cos(2u) = cos^2(u)-sin^2(u)
etc.
-----
tan(2u) = [2tan(u)]/[1-tan^2(u)]
=========
Cheers,
Stan H.
==============